OpenStudy (hockeychick23):

Your interested in predicting college grades (Y) from SAT scores (X). You know that the mean of the SAT scores is 485 with a standard deviation of 85. The average college GPA is 3.1 with a standard deviation of .4. The correlation between SAT scores and college grades is .71. What is the equation of the regression of grades (Y) on scores (X)?

OpenStudy (hockeychick23):

@rational can you check my answer please? (.71)(0.4/85)=0.003 3.1= b+(0.003)(485) b= 1.645 y = 1.645 + .003x

OpenStudy (hockeychick23):

@jim_thompson5910 can you check my answer please?

jimthompson5910 (jim_thompson5910):

one moment

OpenStudy (hockeychick23):

oh i got it, thanks!